Derive a formula for the maximum speed vmax of a simple pendulum bob in terms of g, the length ℓ, and the maximum angle of swing θmax. NOTE!! When entering a trig function of an angle, use parentheses around the argument of the trig function. For instance, enter cos( θmax) rather than cos θmax.
Answer #1

when the pendulum bob is at rightmost position making angle θmax with the vertical

in triangle OBC

Cosθmax = OC/OB

Cosθmax = OC/L

OC = L Cosθmax

height gained relative to the lowest position is given as

h = OA - OC = L - L cos\theta

Using conservation of energy

Potential energy = kinetic energy

(0.5) m V2 = mgh

V2 = = 2g(L - L cos\theta)

Vmax = sqrt (2g(L - L cos\theta))